OpenStudy (anonymous):

how to slove y over x times y to the second times 5y over x ? ( to the simplest from )

OpenStudy (anonymous):

\[5y^4/x^2\]

OpenStudy (anonymous):

but why ?

OpenStudy (anonymous):

y/x * y squared = y^3/x

OpenStudy (anonymous):

Then y^3/x * 5y/x = 5y^4/x^2

OpenStudy (anonymous):

ok .. thx !!

OpenStudy (anonymous):

it's an application of exponant rules

OpenStudy (anonymous):

Multiplying x * x = x^2